given the following solution of potassium permanganate: 

C=2 mol/L

V= 500 ml

a)calculate the volume needed to prepare another solution of potassium permanganate of concentration C=0.2mol/L and V=100mL.

b)50 mL of the prepared solution are taken and 150 mL of water are added. Calculate the new concentration of the solution.


1
Expert's answer
2021-01-11T03:51:05-0500

a) V1 = C2*V2/C1 = 0.2*100/2 = 10 ml

b) C2 = (C1*V1)/(V1+Vwater) = (2*50)/(50+150) = 0.5 mol/L